# Curious QNEIs from QNEC: New Bounds on Null Energy in Quantum Field Theory

### v1

@@ 2.2.2 Bounds on first derivatives of entropy
We will derive bounds between $\partial_{v}S(v)$ using strong subadditivity and monotonicity of relative entropy. [^3] As an aside, note that SSA and monotonicity of relative entropy are equivalent lieb1973proof , which suggests there exists a possible simplification of our following construction. The bound will be in terms of the $\partial_{v}$ derivative of vacuum-subtracted modular Hamiltonians and vacuum entropies.
@@
where, recall, $B(v)$ and $C(v)$ are arbitrary intervals, except that each shares a boundary with $A(v)$, $D(B)\subset D(A)$, and $D(C)\subseteq D(A^{c})$. The modular Hamiltonian and the vacuum entanglement entropy for generic 2d QFTs and arbitrary intervals are not known. However, for a given 2d QFT, if we assume that the null interval limit is controlled by the UV fixed-point CFT of the QFT, then we can use the known CFT results for $\Delta K$ and $S_{\Omega}$ Holzhey:1994we ; Casini:2011kv . We will utilise this in what follows. To take the null interval limit of $B(v)$ and $C(v)$, we set $u_{B}(v)\to 0$ in (25), and $u_{C}(v)\to 0$ in (32).
@@ 2.2.3 Modular energies of null intervals
As mentioned earlier, free theories in $d>2$ display ultralocality when quantized on a lightsheet and the theory decomposes into 2d CFTs along each pencil Wall:2011hj , to each of which we can apply the inversion $v\to v^{\prime}=v^{-1}$. So, (37) (with an additional integration over $\int\text{d}^{d-2}y_{\perp}$) in $d>2$ is correct for free theories, but not expected to be true for interacting theories. In Section 3, we will return to the modular Hamiltonian of a null strip in interacting theories, which can be calculated using that $\Delta K=\Delta S$ Koeller:2015qmn .
@@ 2.2.5 QNEIs
If we compare this to the FH bound (15), we see that, for $a\ll 1$ but finite, this bound is of a similar form, but it is a weaker bound. Since the FH bound is tight, at least for 2d CFTs, we expect that the bounds that we have derived are, in general, weaker than the FH bound, however it unclear if this strictly so or whether by a clever choice of $\zeta$ if we can match the FH bound.
@@ 3 Higher-dimensional quantum null energy inequalities
In this section, we will generalize our two-dimensional approach to derive new QNEIs in $d>2$ dimensions. First, we will introduce the higher-dimensional QNEC, and explain why there is no bound on $\langle T_{vv}\rangle$ smeared along a single null geodesic in higher dimensions, including isolating the point where the proof method used in Sec. 2.2 trivialises the bound. Then, by modifying the method of Sec. 2.2 applied to nearly null infinite strip regions, we will introduce and derive an infinite family of QEIs that involve smearing $\langle T_{vv}(v)\rangle$ along two null directions. This is entirely akin to the “double smeared null energy condition” (DSNEC) introduced in Fliss:2021gdz as a regularization of the SNEC. More specifically, the DSNEC takes the schematic form
$\int\text{d}u\,\text{d}v~g(u,v)\langle T_{vv}\rangle_{\psi}\geq-\frac{\mathcal{N}[\mu]}{\delta_{v}^{\frac{d+2}{2}}\delta_{u}^{\frac{d-2}{2}}}~,\qquad\text{(DSNEC)}$ (63)
where $\delta_{u,v}$ are the smearing length scales of $g$ in the $u$ and $v$ directions, respectively, and $\mathcal{N}[\mu]$ is a function of any correlation lengths $\xi$, measured in units of the invariant smearing length, $\mu^{2}=\delta_{v}\delta_{u}\xi^{-2}$, and scaling with the number of degrees of freedom. In the small $\delta_{u}$ limit with $\delta_{v}$ fixed, this provides a geometrically regulated SNEC
$\int\text{d}u\,\text{d}v~g(u,v)\langle T_{vv}\rangle_{\psi}\overset{\delta_{u}\ll 1}{\geq}-\frac{\mathcal{N}[0]}{\ell_{\text{UV}}^{d-2}}\frac{1}{\delta_{v}^{2}}\left(1+O(\delta_{u})\right)~,$ (64)
where $\ell_{\text{UV}}^{2}=\delta_{u}\delta_{v}$. The DSNEC has been proven for free field theories Fliss:2021phs . Our bounds will be reminiscent of, yet logically distinct from, the DSNEC and, importantly, apply more broadly to generic interacting CFTs.
@@ 3.1 QNEC in higher dimensions
with $\text{d}y_{\perp}^{2}=\sum_{i=1}^{d-2}\text{d}y_{i}^{2}$, and a one-parameter family of regions whose entanglement cuts intersect a constant $u$ null plane at $v=V_{\lambda}(y_{\perp})=V_{0}(y_{\perp})+\lambda\dot{V}(y_{\perp})$. This gives us a one-parameter family of entropies, $S(\lambda)$. Then the QNEC in its general non-local form is  Bousso:2015mna ; Casini:2022rlv :
@@
$\frac{d^{2}S}{d\lambda^{2}}$ is an ordinary derivative. It can be written in terms of functional derivatives as follow:
@@
The off-diagonal piece in (68) is the so-called entanglement density and is non-positive by strong subadditivity Faulkner:2015csl . We can discard the off-diagonal part of (68) by taking a delta function localised shape deformation $\dot{V}(y_{\perp})^{2}=\delta^{(d-2)}(y_{\perp})$. For such a localised shape deformation, QNEC (66) reduces to its local form:
@@
This local form of the QNEC is saturated in interacting CFTs with a twist gap Leichenauer:2018obf ; Balakrishnan:2019gxl .
Next, let us return to QNEC in its general form, eq. (66), and consider a family of null-translated flat entanglement cuts: $V_{0}(y_{\perp})=0$ and $\dot{V}(y_{\perp})=1$. Each entanglement cut in the family is uniform in the transverse directions and parameterized by $\lambda=v$. Then we can integrate (66) against $g(v)$, and integrate by parts, to get
@@
From (70) we can see that the triviality of any putative QNEI must follow from a divergence appearing in the derivative of the entanglement entropy associated with a null interval. We can reinterpret the conclusions of Fliss:2021gdz for free theories from this perspective as well. For instance, consider the free bosonic theory quantized on a lightsheet $\mathcal{L}$. In the procedure we developed in Section 2.2, our lower bounds were set by derivatives of the entanglement entropy reduced on a null interval, and so we consider here a null strip with $v\in[v_{1},v_{2}]$. As emphasized in Wall:2011hj , the vacuum factorizes into the vacuum along a coarse-grained null ray, or “pencil” $\mathfrak{p}$, as
@@ 3.2 A family of interacting QNEIs
Even though our original family of QNEIs in Section 2.2 are (generally) trivial for higher-dimensional theories due to the divergence of the vacuum entanglement entropy, we can regulate this divergence geometrically by smearing over the other null direction as proposed in the DSNEC papers Fliss:2021gdz ; Fliss:2021phs ; however, we will need the further technical ingredient of integrating over the transverse directions. We will return to and leverage the general approach of Section 2.2. Namely, we will bound the stress tensor in a state, smeared over a function $g(u,v)$ with compact support [^9] This assumption of compact support is for the ease of illustration. There should be no difficulty in extending our arguments to Schwartz functions as well. in a causal diamond extended in the $\vec{y}_{\perp}\in\mathbb{R}^{d-2}$ transverse space, in terms of null derivatives of an entanglement entropy through the integrated QNEC, (70):
$\int\text{d}^{d-2}y_{\perp}\int\text{d}u\,\text{d}v\,g(u,v)\langle T_{vv}(u,v,y_{\perp})\rangle$ $\geq\frac{1}{2\pi}\int\text{d}u\,\text{d}v\,g(u,v)\partial_{v}^{2}S(A)$
$=-\frac{1}{2\pi}\int\text{d}u\,\text{d}v\,\partial_{v}g(u,v)\partial_{v}S(A)~.$ (75)
Here $S(A)$ is the entanglement entropy of the state $\psi$ reduced on a region $A$, which is any achronal codimension-1 half-space whose boundary is a flat cut at $(u,v)$ for all $\vec{y}_{\perp}\in\mathbb{R}^{d-2}$:
$\partial A=\Big\{(u,v,y_{\perp})\Big|y_{\perp}\in\mathbb{R}^{d-2}\Big\}~.$ (76)
Note that although we have not notated it explicitly, the region $A$ depends on the point $(u,v)$ in the $u$-$v$ plane and there is a different region $A$ for each point in the integral appearing in (3.2). The difference between this set-up and that of Section 2.2 is that in bounding the first derivatives of $S(A)$, to regulate the UV divergences, we will take our auxiliary system to not be exactly null, but nearly null.
@@
and whose domain of dependence is contained in that of $A$: $\mathcal{D}(B)\subset\mathcal{D}(A)$. Such a region is illustrated on the left of Fig. 3. Much like the region $A$, $B$ must depend explicitly on the point $(u,v)$ in the $u$-$v$ plane, at the very least, by virtue of sharing a common boundary with $A$; we can also use this freedom to choose the second boundary at $\{(u_{B},v_{B},\vec{y}_{\perp})\}$ to vary as we vary $(u,v)$, i.e. $(u_{B},v_{B})\rightarrow(u_{B}(u,v),v_{B}(u,v))$. However in order to maintain $\mathcal{D}(B)\subset\mathcal{D}(A)$ we must have $u_{B}(u,v)>u$ and $v_{B}(u,v)<v$ for all $(u,v)$ of interest. We will see major simplifications in the limit that $B$ is taken to be nearly-null:
$\epsilon_{B}:=u_{B}-u\ll 1~.$ (78)
For the lower bound we will consider a region $C$ to be a nearly null strip
@@
Again $C$ depends on $(u,v)$ by sharing a boundary with $A$ and we can additionally allow its right boundary to vary with $(u,v)$ as well. However we will maintain $u_{C}(u,v)<u$ and $v_{C}(u,v)>v$ such that domain of dependence of $C$ is contained in the domain of the complement of $A$: $\mathcal{D}(C)\subset\mathcal{D}(A^{c})$, as depicted in the right of Figure 3. Again, we will want to take a limit that $C$ becomes null, which will require
$\epsilon_{C}:=u-u_{C}\ll 1$ (80)
for all $(u,v)$.
@@
Figure 3: The regions $A$, $B$, and $C$ in the setup for the higher dimensional derivation. The turquoise region is the domain of support of $g(u,v)$ in the $(u,v)$ plane and is extended uniformly in the remaining $\mathbb{R}^{d-2}$ directions.
@@
As in $d=2$ (see (27) and (34)), $\partial_{v}^{\scriptscriptstyle{(\downarrow)}}$ and $\partial_{v}^{\scriptscriptstyle{(\uparrow)}}$ denote that $\partial_{v}$ acts only on the boundary shared with $A$. In this two-sided bound, all that has changed with respect to $d=2$ is that $A$, $B$ and $C$ have an extension into the transverse space $y_{\perp}\in\mathbb{R}^{d-2}$. But, in $d>2$, the derivatives of the vacuum entropies are not UV-finite when we take the null limit of the strip regions, so we will regulate by taking $B$ and $C$ to be nearly null. However, when $B$ and $C$ are not strictly null, we lose some of the universality of the results of Section 2.2. In particular, we do not have a closed form expression for entanglement entropies of $S(B)$ and $S(C)$, nor local expressions for the vacuum-subtracted modular Hamiltonian. Nonetheless, we can work perturbatively in the nearly-null limit. In particular we can still express
@@
To see why this is so, we evaluate the entanglement entropy through the replica trick. Assuming that our state in question can be prepared by path-integral preparation with operator insertions [^10] By the state operator correspondence such states are dense in a CFT. we can express the entanglement entropy as
@@
where $Z_{n}$ the CFT path-integral on the $n$-fold branched cover, $M_{n}$, with the insertions preparing $\psi$ replicated on each branch. We can map this back to a partition function of a symmetric product CFT, which we denote as $(\text{CFT})^{n}/\mathbb{Z}_{n}$, on a trivial topology. The modding by $\mathbb{Z}_{n}$ means we only allow replica symmetric operators in the spectrum. [^11] As emphasized in Balakrishnan:2019gxl , this projection is necessary for the defect CFT living on the entanglement cut to be a local CFT. The replica trick is encoded by the insertion of twist-1 codim-2 defect operators $D_{n}$ at the boundaries of the strip:
$S(B)=\lim_{n\rightarrow 1}\frac{1}{1-n}\log\langle D_{n}(u_{B},v_{B})D_{n}(u,v)\rangle~.$ (84)
The twist defect operators admit an operator product expansion (OPE) of the following form [^12] The coefficient of (1n)(1-n) in the exponential is assumed based on the assumption of analyticity of the Rényi entropy in $n$ Headrick:2010zt . Bousso:2014uxa
@@
The lightcone limit $u\rightarrow 0$ of (85) is very similar to a coincident limit of this OPE and takes a simplified form Casini:2017roe
@@
where we have expressed the replica symmetric operator explicitly as a sum over seed stress tensors and indicated schematically the sum over its descendants. We will make the additional assumption that we can expand this OPE about $n=1$ to arrive at
@@
Because the VEV of the single-copy stress-tensors vanishes, the identity term of (3.2) is the leading contribution to the vacuum entanglement entropy $S_{\Omega}(B)$. The stress tensor contribution instead determines the entropy difference $\Delta S$. Note that this second term is finite in the null limit. In the strict null limit, $u\rightarrow 0$, this is also equal to the vacuum subtracted modular Hamiltonian, $\Delta S=\Delta K$, due to the vanishing of the relative entropy for interacting theories on a null sheet Bousso:2014uxa . From this finiteness we will assume that it admits an expansion in the separation in $u$-direction
$\Delta S(B)=2\pi\int d^{d-2}y_{\perp}\int_{v_{B}}^{v}\text{d}v^{\prime}(v-v_{B})G\left(\frac{v^{\prime}-v_{B}}{v-v_{B}}\right)\langle T_{vv}(u,v^{\prime},y_{\perp})\rangle+O(\epsilon_{B})~,$ (90)
Here $G$ encapsulates the resummation of the descendants in the stress tensor channel of the twist defect OPE. Thus it is theory specific. For instance, the free theory modular Hamiltonian is given by G(s)=s(1s)G(s)=s(1-s) (consistent with what we have seen in Section 2.2), however is known to differ for holographic CFTs Bousso:2014uxa . Regardless, $G(s)$ obeys several properties that were derived in Bousso:2014uxa :
@@
$\partial_{v}^{\scriptscriptstyle{(\downarrow)}}\Delta S(B)=2\pi\int d^{d-2}y_{\perp}\int_{v_{B}}^{v}\text{d}v^{\prime}\mathcal{G}\left(\frac{v^{\prime}-v_{B}}{v-v_{B}}\right)\langle T_{vv}(u,v^{\prime},y_{\perp})\rangle+O(\epsilon_{B})~,$ (93)
@@
We now focus on the divergent and finite contributions to the vacuum entropy. It is natural to assume, as we have above, that the coefficient for the identity operator term in the defect OPE is proportional to the prefactor $c_{\scriptscriptstyle\mathrm{T}}$ of the stress-tensor two-point function [^13] We use the normalization convention of Osborn:1993cr , which is most easily stated in a time-like colinear frame as $\langle T_{00}(t)T_{00}(0)\rangle_{\Omega}=\frac{(d-1)}{d}\frac{c_{\scriptscriptstyle\mathrm{T}}}{t^{2d}}~.$ (95) . This assumption can be corroborated through the calculation of a nearly null strip in a holographic CFT using the Ryu-Takayanagi formula Ryu:2006bv . In this context, the bulk Newton’s constant is proportional to $c_{\scriptscriptstyle\mathrm{T}}$ Buchel:2009sk which fixes the coefficient $\beta_{d}$ to be
@@
although we acknowledge that this coefficient may differ for non-holographic CFTs; in any case, its precise form will not be important for what follows. In the $\epsilon_{B}=(u_{B}-u)\rightarrow 0$ limit the defect OPE computation suggests that the vacuum entanglement entropy of a strip in the null limit takes a universal form in interacting CFTs [^14] It is feasible that a generic interacting QFT might also have significant simplifications in its OPE in the null limit. However because we are using features of the defect OPE (i.e. convergence of the OPE) that are have only been strictly established for CFTs we will be cautious and specify CFTs in this section.:

### v2

@@ 2.2.2 Bounds on first derivatives of entropy
We will derive bounds between $\partial_{v}S(v)$ using strong subadditivity and monotonicity of relative entropy. [^3] As an aside, note that SSA and monotonicity of relative entropy are equivalent Lieb and Ruskai (1973), which suggests there exists a possible simplification of our following construction. The bound will be in terms of the $\partial_{v}$ derivative of vacuum-subtracted modular Hamiltonians and vacuum entropies.
@@
where, recall, $B(v)$ and $C(v)$ are arbitrary intervals, except that each shares a boundary with $A(v)$, $D(B)\subset D(A)$, and $D(C)\subseteq D(A^{c})$. The modular Hamiltonian and the vacuum entanglement entropy for generic 2d QFTs and arbitrary intervals are not known. However, for a given 2d QFT, if we assume that the null interval limit is controlled by the UV fixed-point CFT of the QFT, then we can use the known CFT results for $\Delta K$ and $S_{\Omega}$ Holzhey et al. (1994); Casini et al. (2011). We will utilise this in what follows. To take the null interval limit of $B(v)$ and $C(v)$, we set $u_{B}(v)\to 0$ in (25), and $u_{C}(v)\to 0$ in (32).
@@ 2.2.3 Modular energies of null intervals
As mentioned earlier, free theories in $d>2$ display ultralocality when quantized on a lightsheet and the theory decomposes into 2d CFTs along each pencil Wall (2012a), to each of which we can apply the inversion $v\to v^{\prime}=v^{-1}$. So, (37) (with an additional integration over $\int\text{d}^{d-2}y_{\perp}$) in $d>2$ is correct for free theories, but not expected to be true for interacting theories. In Section 3, we will return to the modular Hamiltonian of a null strip in interacting theories, which can be calculated using that $\Delta K=\Delta S$ Koeller and Leichenauer (2016).
@@ 2.2.5 QNEIs
If we compare this to the FH bound (15), we see that, for $a\ll 1$ but finite, this bound is of a similar form, but it is a weaker bound. Since the FH bound is tight, at least for 2d CFTs, we expect that the bounds that we have derived are, in general, weaker than the FH bound. However, it is unclear if this is strictly so or whether, by a clever choice of $\zeta$, we can match the FH bound.
@@ 3 Higher-dimensional quantum null energy inequalities
In this section, we will generalize our two-dimensional approach to derive new QNEIs in $d>2$ dimensions.
@@ 3.1 QNEC in higher dimensions
with $\text{d}y_{\perp}^{2}=\sum_{i=1}^{d-2}\text{d}y_{i}^{2}$, and a one-parameter family of regions whose entanglement cuts intersect a constant $u$ null plane at $v=V_{\lambda}(y_{\perp})=V_{0}(y_{\perp})+\lambda\dot{V}(y_{\perp})$. This gives us a one-parameter family of entropies, $S(\lambda)$. Then the QNEC in its general non-local form is  Bousso et al. (2016b); Casini and Huerta (2023):
@@
$\frac{d^{2}S}{d\lambda^{2}}$ is an ordinary derivative. It can be written in terms of functional derivatives as follows:
@@
The off-diagonal piece in (66) is the so-called entanglement density and is non-positive by strong subadditivity Faulkner et al. (2016b). We can discard the off-diagonal part of (66) by taking a delta function localised shape deformation $\dot{V}(y_{\perp})^{2}=\delta^{(d-2)}(y_{\perp})$. For such a localised shape deformation, QNEC (64) reduces to its local form:
@@
This local form of the QNEC is saturated in interacting CFTs with a twist gap Leichenauer et al. (2018); Balakrishnan et al. (2022).
Next, let us return to QNEC in its general form, eq. (64), and consider a family of null-translated flat entanglement cuts: $V_{0}(y_{\perp})=0$ and $\dot{V}(y_{\perp})=1$. Each entanglement cut in the family is uniform in the transverse directions and parameterised by $\lambda=v$. Then we can integrate (64) against $g(v)$, and integrate by parts, to get
@@
At this point, we could continue to follow the steps of Section 2.2 to attempt to derive a higher-dimensional QNEI. But the attempt is doomed: no such bound exists. We have already hinted at this in our discussion at the end of that section, namely that within the context of free field theories, it is possible to construct a series of squeezed states with no lower bound on their null-energy density integrated along a single null geodesic. Also, as pointed out in Fliss and Freivogel (2022), further integrating this energy density along the transverse directions does not change this conclusion. In addition to the concrete counterexamples in free theory, there exists a general argument against a state-independent lower bound on null energy density integrated over a finite segment of a null geodesic which goes as follows.
@@
$T_{vv}(\lambda u,\lambda v,\lambda\vec{y}_{\perp})=\lambda^{-d}U_{D(\lambda)}T_{vv}(u,v,y_{\perp})U^{\dagger}_{D(\lambda)}~.$ (69)
Similarly, under boosts there exists a unitary operator $U_{K(\lambda)}$ such that
$T_{vv}(\lambda u,\lambda^{-1}v,\vec{y}_{\perp})=\lambda^{2}U_{K(\lambda)}T_{vv}(u,v,y_{\perp})U^{\dagger}_{K(\lambda)}~.$ (70)
[…]
$\int\text{d}v\,g(v)\,\langle T_{vv}(0,v,\vec{0})\rangle_{\psi}=-\mathcal{F}~,$ (71)
[…]
$\int\text{d}v\,g(v)\langle T_{vv}(0,v,\vec{0})\rangle_{\tilde{\psi}}=-\lambda^{d-2}\mathcal{F}~.$ (72)
Then $\int\text{d}v\,g(v)\,\langle T_{vv}\rangle$ can be made arbitrarily negative and so cannot admit a state-independent lower bound.
From (68) we can alternatively track the triviality of any putative QNEI from a divergence appearing in the derivative of the entanglement entropy associated with a null interval. We can reinterpret the conclusions of Fliss and Freivogel (2022) for free theories from this perspective as well. For instance, consider the free bosonic theory quantized on a lightsheet $\mathcal{L}$. In the procedure we developed in Section 2.2, our lower bounds were set by derivatives of the entanglement entropy reduced on a null interval, and so we consider here a null strip with $v\in[v_{1},v_{2}]$. As emphasized in Wall (2012a), the vacuum factorizes into the vacuum along a coarse-grained null ray, or “pencil” $\mathfrak{p}$, as
@@ 3.2 A family of interacting QNEIs
We will return to and leverage the general approach of Section 2.2. Namely, we will bound the stress tensor in a state, smeared over a function $g(v)$ with compact support [^8] This assumption of compact support is for the ease of illustration. There should be no difficulty in extending our arguments to Schwartz functions as well. in a causal diamond extended in the $\vec{y}_{\perp}\in\mathbb{R}^{d-2}$ transverse space, in terms of null derivatives of an entanglement entropy through the integrated QNEC, (68):
$\int\text{d}^{d-2}y_{\perp}\int\,\text{d}v\,g(v)\langle T_{vv}(u_{0},v,y_{\perp})\rangle$ $\geq\frac{1}{2\pi}\int\text{d}u\,\text{d}v\,g(v)\partial_{v}^{2}S(A)$
$=-\frac{1}{2\pi}\int\text{d}v\,\partial_{v}g(v)\partial_{v}S(A)~.$ (77)
Here $S(A)$ is the entanglement entropy of the state $\psi$ reduced on a region $A$, which is any achronal codimension-1 half-space whose boundary $\partial A$ is a flat cut at $(u_{0},v)$ for all $\vec{y}_{\perp}\in\mathbb{R}^{d-2}$:
$\partial A(v)=\Big\{(u_{0},v,y_{\perp})\Big|y_{\perp}\in\mathbb{R}^{d-2}\Big\}~.$ (78)
The region $A$ depends on the value of $v$, and there is a different region for each point in the integral appearing in (3.2). The difference between this set-up and that of Section 2.2 is that in bounding the first derivatives of $S(A)$, to regulate the UV divergences, we will take our strip regions to not be exactly null, but nearly null.
@@
and whose domain of dependence is contained in that of $A$: $\mathcal{D}(B)\subset\mathcal{D}(A)$. Such a region is illustrated on the left of Fig. 3. Much like the region $A$, $B$ must depend explicitly on the point $(u_{0},v)$ in the $u$-$v$ plane, at the very least, by virtue of sharing a common boundary with $A$; we can also use this freedom to choose the second boundary at $\{(u_{B},v_{B},\vec{y}_{\perp})\}$ to vary as we vary $v$, i.e. $(u_{B},v_{B})\rightarrow(u_{B}(v),v_{B}(v))$. However in order to maintain $\mathcal{D}(B)\subset\mathcal{D}(A)$ we must have $u_{B}(v)>u_{0}$ and $v_{B}(u,v)<v$ for all $v$ of interest. We will see major simplifications in the limit that $B$ is taken to be nearly null:
$\epsilon_{B}:=u_{B}-u_{0}\ll 1~.$ (80)
For the lower bound, we will consider a region $C$ to be a nearly null strip
@@
Again, $C$ depends on $v$ by sharing a boundary with $A$, and we can additionally allow its right boundary to vary with $v$ as well. However we will maintain $u_{C}(v)<u_{0}$ and $v_{C}(v)>v$ such that domain of dependence of $C$ is contained in the domain of the complement of $A$: $\mathcal{D}(C)\subset\mathcal{D}(A^{c})$, as depicted in the right of Figure 3. Again, we will want to take a limit that $C$ becomes null, which will require
$\epsilon_{C}:=u_{0}-u_{C}\ll 1$ (82)
for all $v$.
@@
Figure 3: The $v$-dependent family of codim-11 regions $A$, $B$, and $C$ in the setup for the higher dimensional derivation, all of which extend uniformly in the transverse $\mathbb{R}^{d-2}$ directions. AA is a half-plane and $B$ and $C$ are strips.
@@
As in $d=2$ (see (27) and (34)), $\partial_{v}^{\scriptscriptstyle{(\downarrow)}}$ and $\partial_{v}^{\scriptscriptstyle{(\uparrow)}}$ denote that $\partial_{v}$ acts only on the boundary shared with $A$. In this two-sided bound, all that has changed with respect to $d=2$ is that $A$, $B$ and $C$ have an extension into the transverse space $y_{\perp}\in\mathbb{R}^{d-2}$. But, in $d>2$, the derivatives of the vacuum entropies are not UV-finite when we take the null limit of the strip regions, so we will regulate by taking $B$ and $C$ to be nearly null. However, when $B$ and $C$ are not strictly null, we lose some of the universality of the results of Section 2.2. In particular, we do not have a closed-form expression for entanglement entropies of $S(B)$ and $S(C)$, nor local expressions for the vacuum-subtracted modular Hamiltonian. Nonetheless, we can work perturbatively in the nearly-null limit. In particular, we can still express
@@
To see why this is so, we evaluate the entanglement entropy through the replica trick. Assuming that our state in question can be prepared by path-integral preparation with operator insertions [^9] By the state operator correspondence, such states are dense in a CFT. we can express the entanglement entropy as
@@
where $Z_{n}$ the CFT path-integral on the $n$-fold branched cover, $M_{n}$, with the insertions preparing $\psi$ replicated on each branch. We can map this back to a partition function of a symmetric product CFT, which we denote as $(\text{CFT})^{n}/\mathbb{Z}_{n}$, on a trivial topology. The modding by $\mathbb{Z}_{n}$ means we only allow replica symmetric operators in the spectrum. [^10] As emphasized in Balakrishnan et al. (2022), this projection is necessary for the defect CFT living on the entanglement cut to be a local CFT. The replica trick is encoded by the insertion of twist-1 codim-2 defect operators $D_{n}$ at the boundaries of the strip:
$S(B)=\lim_{n\rightarrow 1}\frac{1}{1-n}\log\langle D_{n}(u_{B},v_{B})D_{n}(u_{0},v)\rangle~.$ (86)
The twist defect operators admit an operator product expansion (OPE) of the following form [^11] The coefficient of (1n)(1-n) in the exponential is assumed based on the assumption of analyticity of the Rényi entropy in $n$ Headrick (2010). Bousso et al. (2015)
@@
The lightcone limit $u\rightarrow 0$ of (87) is very similar to a coincident limit of this OPE and takes a simplified form Casini et al. (2017)
@@
where we have expressed the replica symmetric operator explicitly as a sum over seed stress tensors and indicated the sum over its descendants schematically. We will make the additional assumption that we can expand this OPE about $n=1$ to arrive at
@@
Because the VEV of the single-copy stress-tensors vanishes, the identity term of (3.2) is the leading contribution to the vacuum entanglement entropy $S_{\Omega}(B)$. The stress tensor contribution instead determines the entropy difference $\Delta S$. Note that this second term is finite in the null limit. In the strict null limit, $u\rightarrow 0$, this is also equal to the vacuum subtracted modular Hamiltonian, $\Delta S=\Delta K$, due to the vanishing of the relative entropy for interacting theories on a null sheet Bousso et al. (2015). From this finiteness, we will assume that it admits an expansion in the separation in the $u$-direction
$\Delta S(B)=2\pi\int d^{d-2}y_{\perp}\int_{v_{B}}^{v}\text{d}v^{\prime}(v-v_{B})G\left(\frac{v^{\prime}-v_{B}}{v-v_{B}}\right)\langle T_{vv}(u_{0},v^{\prime},y_{\perp})\rangle+O(\epsilon_{B})~,$ (92)
Here $G$ encapsulates the resummation of the descendants in the stress tensor channel of the twist defect OPE. Thus it is theory-specific. For instance, the free theory modular Hamiltonian is given by G(s)=s(1s)G(s)=s(1-s) (consistent with what we have seen in Section 2.2), however is known to differ for holographic CFTs Bousso et al. (2015). Regardless, $G(s)$ obeys several properties that were derived in Bousso et al. (2015):
@@
$\partial_{v}^{\scriptscriptstyle{(\downarrow)}}\Delta S(B)=2\pi\int d^{d-2}y_{\perp}\int_{v_{B}}^{v}\text{d}v^{\prime}\mathcal{G}\left(\frac{v^{\prime}-v_{B}}{v-v_{B}}\right)\langle T_{vv}(u_{0},v^{\prime},y_{\perp})\rangle+O(\epsilon_{B})~,$ (95)
@@
We now focus on the divergent and finite contributions to the vacuum entropy. It is natural to assume, as we have above, that the coefficient for the identity operator term in the defect OPE is proportional to the prefactor $c_{\scriptscriptstyle\mathrm{T}}$ of the stress-tensor two-point function [^12] We use the normalization convention of Osborn and Petkou (1994), which is most easily stated in a time-like colinear frame as $\langle T_{00}(t)T_{00}(0)\rangle_{\Omega}=\frac{(d-1)}{d}\frac{c_{\scriptscriptstyle\mathrm{T}}}{t^{2d}}~.$ (97) . This assumption can be corroborated through the calculation of a nearly null strip in a holographic CFT using the Ryu-Takayanagi formula Ryu and Takayanagi (2006). In this context, the bulk Newton’s constant is proportional to $c_{\scriptscriptstyle\mathrm{T}}$ Buchel et al. (2010) which fixes the coefficient $\beta_{d}$ to be
@@
although we acknowledge that this coefficient may differ for non-holographic CFTs; in any case, its precise form will not be important for what follows. In the $\epsilon_{B}=(u_{B}-u_{0})\rightarrow 0$ limit, the defect OPE computation suggests that the vacuum entanglement entropy of a strip in the null limit takes a universal form in interacting CFTs [^13] It is feasible that a generic interacting QFT might also have significant simplifications in its OPE in the null limit. However, because we are using features of the defect OPE (i.e. convergence of the OPE) that have only been strictly established for CFTs, we will be cautious and specify CFTs in this section.: